The psfonts package^{∗}

Sebastian Rahtz S.Rahtz@elsevier.co.uk

1995/04/01

1 Introduction

This file contains suitable package files to load the standard PostScript fonts. The font definition files and font metrics are available in both T1 and OT1 form in the companion CTAN fonts/metrics collection.

∗This file has version number 5.0, last revised 1995/04/01.